One number is 17 more than another. If the greater number is increased by 3, the result is 5 times the smaller. What are the numbers? I have had everyone I know help me with this and we cannot work this problem. My last attempt yield me 17.75 greater number and 3.75 small.........that does not work. Help!!
Answer by ptaylor(2198)

Let x=the smaller number; 5x=five times the smaller number
Then x+17=the larger number
x+17+3=x+20=larger number increased by 3
Now we are told that:
x+20=5x subtract x from each side
x-x+20=5x-x collect like terms
20=4x divide each side by 4
x=5-----------------------the smaller number
x+17=5+17=22 -------------the larger number
CK
22+3=5*5
25=25
Hope this helps---ptaylor